OpenStudy (anonymous):

x^2-1/3x What number should be added to make this a perfect square trinomial? help me solve pls

OpenStudy (tkhunny):

Try it out! (a+b)^2 = a^2 + 2ab + b^2 Q: How does the second term realte to teh third term? A: You have "2b". How can you find "b^2"?

OpenStudy (anonymous):

well I got the answer as it was at the bottom of the practice if you do my problem what answer do you come up with? it should be 1/36 I just can't figure out how to get it

OpenStudy (tkhunny):

You have 2b. In your problem statement, that is -1/3. Calcualte b. In your problem statement, that is -1/6 (dividing by 2) Calculate b^2. In your problem statement, that is (-1/6)^2 = 1/36